# MeteoRA: Multiple-tasks Embedded LoRA for Large Language Models

### v1

@@ Appendix A Appendix
Table 5, Table 6, Table 7 and Table 8 show the detailed evaluation results of different models on the 28 selected tasks.
@@
| abstract_narrative_understanding | 0.803 | 0.787 | 0.787 | 0.793 | 0.79 |
@@
| topical_chat | PEFT (reference) | 35.60 | 0.150 | 0.024 | 0.140 |
@@
A.3 composite-n evaluation results details
Table 9, Table 10 and Table 11 are detailed evaluation results for composite-3, composite-5 and composite-10 respectively. There are a few points that need additional clarification here. (1) The models are required to generate both the corresponding question number and its answer. Thus, any mismatch between the question number and answer is considered incorrect when evaluating. (2) In the evaluation results below, some BLEU scores are because the model generated mismatched question numbers and answers or extremely brief answers, leading to an overall BLEU score. (3) The task strategyqa involves answering a question with either ’yes’ or ’no’ and providing reasoning steps. The accuracy metric for this task refers to the ’yes’ or ’no’ accuracy. (4) The ROUGE scores refer to F1-scores. (5) A sample will be skipped if it is too long.
@@
A.4 Full examples of composite-n task
Below is the full example shown in Figure 4.
@@ A.4 Full examples of composite-n task
logical_deduction question_selection strategyqa

### v2

@@ A.4 Experimental results of 28 tasks
Table 4, Table 5, Table 6 and Table 7 show the detailed evaluation results of different models on the 28 selected tasks. When drawing Figure 3, for tasks we use BLEU and ROUGE as metrics, we selected BLEU for news_commentary_de, news_commentary_es, and news_commentary_it, while opting for ROUGE-L for the remaining tasks.
@@
| abstract_narrative_understanding | 0.803 | 0.787 | 0.787 | 0.793 | 0.790 |
@@
| topical_chat | PEFT (reference) | 12.50 | 0.157 | 0.027 | 0.146 |
@@
A.5 Composite-n evaluation results details
The tasks for constructing composite-n are selected from the aforementioned set of 28 tasks to ensure the models familiarity and potential problem-solving capability. However, given the limited capability of the instruction following in the zero-shot setting, neither the MeteoRA models nor the models fine-tuned by LoRA achieve satisfactory results. Hence, we employ a 2-shot setting for evaluation on these composite-n tasks.
The evaluation metrics used for composite-n tasks are: average number of questions attempted, average number of multiple-choice questions answered correctly, and average BLEU, ROUGE scores for non-multiple-choice questions.
Notice that in the composite-n tasks, when calculating the softmax values of the weights for the two LoRA adapters selected by the Gating network, we introduced a hyperparameter called temperature. The value of temperature needs to be increased as the number of sub-tasks grows. Specifically, we set the temperature values to , , and for the three tasks, respectively.
Tables 8, 9, and 10 present the detailed evaluation results for the composite-3, composite-5, and composite-10 tasks, respectively. Several important clarifications are necessary for interpreting these results:
1.
The models are required to generate both the corresponding question number and its answer. Any mismatch between the question number and the answer is therefore considered incorrect.
2.
In the evaluation results, some BLEU scores are recorded as . This occurs when the model generates mismatched question numbers and answers or provides extremely insufficient answers, resulting in an overall BLEU score.
3.
For the task strategyqa, which involves answering with either ’yes’ or ’no’ and providing reasoning steps, the accuracy metric specifically measures the correctness of the ’yes’ or ’no’ response.
4.
The reported ROUGE scores refer to the F1-scores.
5.
Samples that the lengths exceed to 4,096 tokens are skipped in the evaluation process (we skip 13 samples in total).
@@ A.5 Composite-n evaluation results details
A.6 Full content of the illustarted composite-n task
Below is the full content of a composite-3 task, which includes 2-shot examples, the input (3 problems), and the model generated output, shown in Figure 4.
@@ A.6 Full content of the illustarted composite-n task
logical_deduction question_selection strategyqa other task
A.7 Efficiency evaluation experiments on different MeteoRA forward pass implementations
In addition to experiments on our 28 selected tasks, we assess the efficiency of our MeteoRA forward pass design using randomly-generated pseudo data across various settings, including batch size (), sequence length (), gating weights top-k (), LoRA rank size (), number of LoRAs (), maximum tokens to generate (), input hidden dimension (), and output hidden dimension (). Moreover, here we introduce a new baseline, loop-speedup, which improves upon loop-original by removing redundant or inefficient operations directly, acting like a strong substitute for the original design.
As depicted in Figures 7 and 8, our bmm-torch design outperforms other implementations, boasting an average speedup of over loop-original. However, its memory usage escalates with longer sequence lengths. In contrast, bmm-triton maintains a comparable memory footprint to the baselines while retaining of the speedup achieved by bmm-torch, showcasing a balanced trade-off between time and space, as illustrated in Figure 6.